help solve for v. i got answer from other site but dont understand how they got answer:
3u=5+7v/v-3

the 5+7v is a fraction over v-3

Answer by stanbon(75887)

3u=(5+7v)/(v-3)
-------
Multiply both sides by v-3 to get:
3u(v-3) = 5+7v
--------
3uv - 9u = 5 + 7v
------
Get the v-terms together::
3uv - 7v = 5 + 9u
-----
Factor out the "v":
v(3u-7) = 5+9u
-------
v = (5+9u)/(3u-7)
========================
Cheers,
Stan H.
================